OpenStudy (anonymous):

solve sin(5pi/12); use pi/4+pi/6

OpenStudy (anonymous):

pi/4 + pi/6 = 5pi/12 sin(a+b) = sin(a)*cos(b) + cos(a)*sin(b) so sin(pi/4 + pi/6) =...?

OpenStudy (anonymous):

√2/2 *√3/2 + √2/2 * 1/2 6/4 = 3/2 + √2/4 = 6√2/4 ??

OpenStudy (anonymous):

√2/2 *√3/2 = √6/4

OpenStudy (anonymous):

@Algebraic! so the answer is not √6+√2 / 4?

OpenStudy (anonymous):

that's the right answer.

OpenStudy (anonymous):

thank you